MOA-II Obsevation Log
Return to index
Previous day
Next day
DATE: 2007/08/16
OBSERVER: Ohnishi Paul
CONDITION: Coudy, overcast, snow
COMMENT: Dark D2579(60sec) D2580(60sec) are OK. Peak of MOA373(gb13) is followup. After the midnight. vacuum become worse, so we start the pumping at AM 2:00.
Compressor gas pressure: 1.8 MPa
Expected Observation Time: START 18:59 END 06:44
TARGET | RUN | TOTAL RUN | TOTAL FIELD | TOTAL EXPOSURE (sec) |
GB | 9770-9788 | 19 | 38 | 2280 |
LMC | - | 0 | 0 | 0 |
SMC | - | 0 | 0 | 0 |
ALERT | - | 0 | 0 | 0 |
GRB | - | 0 | 0 | 0 |
TOTAL | --- | 19 | 38 | 2280 (00h38m00s) |
GB FIELD STATISTICS
Field1 | Field2 | Field3 | Field4 | Field5 | Field6 | Field7 | Field8 | Field9 | Field10 | Field11 | Field12 |
1 | 1 | 1 | 1 | 8 | 1 | 0 | 0 | 7 | 0 | 0 | 0 |
Field13 | Field14 | Field15 | Field16 | Field17 | Field18 | Field19 | Field20 | Field21 | Field22 | Field23 |
18 | 0 | 0 | 0 | 0 | 0 | 0 | 0 | 0 | 0 | 0 |
LMC FIELD STATISTICS
Field1 | Field2 | Field3 | Field4 | Field5 | Field6 | Field7 | Field8 | Field9 | Field10 | Field11 | Field12 |
0 | 0 | 0 | 0 | 0 | 0 | 0 | 0 | 0 | 0 | 0 | 0 |
Field13 | Field14 | Field15 | Field16 |
0 | 0 | 0 | 0 |
SMC FIELD STATISTICS
Field1 | Field2 | Field3 | Field4 | Field5 | Field6 | Field7 | Field8 | Field9 | Field10 | Field11 | Field12 |
0 | 0 | 0 | 0 | 0 | 0 | 0 | 0 | 0 | 0 | 0 | 0 |
SPECIAL TARGET NAMEBR>
CHECK SHEET
Time | Status | Weather |
18:00 | waiting | overcast |
19:00 | waiting | cloudy |
20:00 | waiting | overcast |
21:00 | waiting | cloudy |
22:00 | waiting | cloudy |
23:00 | waiting | overcast |
00:00 | waiting | overcast |
01:00 | waiting | overcast |
02:00 | waiting | overcast, Start pumping |
03:00 | waiting | overcast, pumping |
04:00 | waiting | snow, pumping |
05:00 | waiting | snow, pumping |
06:00 | waiting | pumping |
07:00 | waiting | pumping |
CCD TEMPERATURE
Previous day
Next day

VACUUM
Previous day
Next day

TELESCOPE TEMPERATURE
Previous day
Next day

Return to index
Previous day
Next day